# A new understanding of Einstein-Rosen bridges

### v1

@@ 6.2 The mathematical bridges in quantum black hole
Figure 5: The picture represents the spacetime conformal diagram of quantum SBH according to DQFT. It contains four regions I,II,III,IVI,\,II,\,III,\,IV, which define geometric SSS to describe quantum fields in Schwarzschild spacetime (applying the near-horizon approximation $r\approx 2GM$). In this picture, the region II (IIIIII) and IIII (IVIV) are related by discrete transformation $\left(U,\,V\right)\to\left(-U,\,-V\right)$. The curved black lines with arrows represent integral curves of killing vector $\partial_{t}$ (54) in each region or in other words, these are the curves of $T^{2}-X^{2}={\rm constant>0}$ (Region I and $I$) and $T^{2}-X^{2}={\rm constant}<0$ (Region III and IV). It is trivial to see an analogy between the spacetime of Schwarzschild BH and the phase space of IHO (23). The red horizontal lines are identified with $r=0$. The quantum field components in the exterior are $\hat{\Phi}_{I}$ and $\hat{\Phi}_{II}$ at parity conjugate regions with opposite arrows of time $t:$-$\infty\to\infty$ and $t:\infty\to$-$\infty$ respectively. Whereas in the interior quantum field (component) $\hat{\Phi}_{int}$ evolve according to (151).
@@ 6.2.1 A new understanding of ER=EPR
In all the investigations that admit unitarity is lost and usual entanglement between BH interior and exterior states, the basic fact that the concept of ”time” in the interior and exterior is different is unnoticed at the quantum level. It is well known that the radial coordinate $r$ behaves like time and $t$ becomes spacelike in the BH interior, which makes the interior Schwarzschild spacetime not static but analogous to the Kantowski-Sachs cosmological metric [93]. All QFTCS construction requires us to define a positive energy state with respect to an arrow of time, since time is not the same in the interior and exterior of BH; one must have a different Hilbert space structure for these two regions. This is exactly what our concept of geometric SSS does. Here, we present an alternative and complementary formulation of ER=EPR grounded in DQFT. By studying quantum fields across the Rindler and Schwarzschild horizons, we propose that the entanglement structure implied by ER=EPR arises naturally through the direct-sum Hilbert space formalism, without invoking geometric wormholes. To be precise, the outcomes of DQFT (137) and (154) indicate a pure state ($|\Psi_{P}\rangle$) is split into direct-sum of two pure state components $\left(|\Psi_{wH}\rangle,\,|\Psi_{bH}\right)$ that geometrically evolve within and beyond the Horizon as
@@ 6.3.1 Direct-sum QFT in flat FLRW de Sitter
$\hat{\phi}_{\pm}$ $=\int\frac{d^{3}k}{\left(2\pi\right)^{3/2}\sqrt{2k}}\left[d_{(\pm)\,\textbf{k}}\phi_{(\pm)k}\left(\pm\tau\right)e^{\pm i\textbf{k}\cdot\textbf{x}}+d_{(\pm)\,\textbf{k}}\phi_{(\pm)k}^{\ast}\left(\pm\tau\right)e^{\mp i\textbf{k}\cdot\textbf{x}}\right]$
@@ 7 ERBs, direct-sum inflation (DSI) and CMB
This section marks a pivotal point in our exposition, where the theoretical framework developed thus far, including inverted harmonic oscillator dynamics, quantum field theory in spacetimes with horizons, and the direct-sum quantization formalism, culminates in potential observational consequences. Specifically, we propose a novel connection between the mathematical structures underpinning DQFT and anomalies observed in the Cosmic Microwave Background [94], particularly those related to parity asymmetry. Furthermore, we provide a new prediction for primordial gravitational waves for the future observational probes.
The inflationary universe, modelled effectively by quasi-de Sitter spacetime, offers a natural setting where the horizon-induced quantum correlations may leave imprints on primordial fluctuations. In this context, we introduce the idea of Direct-Sum Inflation (DSI): a framework in which quantum states spanning parity conjugated regions, connected via mathematical bridges akin to those in ER=EPR, may manifest observable signatures in the large-scale CMB anisotropies. This opens a concrete avenue for testing the deep theoretical ideas developed throughout this work, providing a possible empirical handle on the quantum structure of spacetime near horizons.
@@ 7.1 Direct-sum QFT of inflationary quantum fluctuations
Inflationary background by definition quasi-dS expansion, and it breaks the symmetry (61) of (59) i.e., $\left(\tau,\,\textbf{x}\right)\to\left(-\tau,\,-\textbf{x}\right)$ by the time-dependent slow-roll parameters $\left(\epsilon,\,\eta\right)$ (82). Recall the fact that time is a parameter in quantum theory (in the sense that time reversal is anti-unitary in character), which we extensively discussed in Sec. 5. In the context of inflationary quantum fluctuations, we literally promote the gravitational and matter (inflaton) field variables as operators around the dynamically evolving quasi-de Sitter background. This is nothing but linearized quantum gravity, as extensively discussed in [95], because we do the following operationally.
@@
with $H^{(1)}_{\nu_{s}}\left(z\right)$ is the Hankel functions of the first kind. The standard inflation (SI) prediction [71] corresponds to $\Delta\mathcal{P}_{v}=0$, while the DSI one is given by Equation (177). Note that the power spectrum is the Fourier transform of the two-point function of curvature perturbation as shown in (176) and (177), which contains two unequal contributions due to $\mathcal{P}\mathcal{T}$ symmetry of dS (58) is broken during inflation. In other words our direct-sum vacuum $|0_{qdS}\rangle=|0_{qdS+}\rangle\oplus|0_{qdS-}\rangle$ is $\mathcal{P}\mathcal{T}$ asymmetric. We can express (178) as the two-point correlation of momentum modes as
@@
$|\Psi_{sq}\rangle=\begin{pmatrix}\prod_{\mathbf{k}}\frac{1}{|\alpha^{+}_{k}|}\exp\!\Big(-\frac{\beta_{k}^{+}}{2\alpha_{k}^{+\ast}}\,\hat{b}^{\dagger}_{\mathbf{k},\,+}\hat{b}_{-\textbf{k},\,+}^{\dagger}\Big)|0_{+}\rangle_{\rm qdS}\\ \prod_{\mathbf{k}}\frac{1}{|\alpha^{-}_{k}|}\exp\!\Big(-\frac{\beta_{k}^{-}}{2\alpha_{k}^{-\ast}}\,\hat{b}^{\dagger}_{\mathbf{k},\,-}\hat{b}_{-\textbf{k},\,-}^{\dagger}\Big)|0_{-}\rangle_{\rm qdS}\end{pmatrix}=\begin{pmatrix}|\Psi_{sq}^{+}\rangle\\ |\Psi_{sq}^{-}\rangle\end{pmatrix}$ (181)
@@
$P_{h}=A_{t}\left(\frac{k}{k_{\ast}}\right)^{n_{t}},\,r=-8n_{t}\,,\quad\Delta\mathcal{P}_{u}=\left(-\frac{r}{8}\right)\operatorname{Re}\left[\frac{2}{H_{3/2}^{(1)}\left(\frac{k}{k_{\ast}}\right)}\frac{\partial H^{(1)}_{\nu_{s}}\left(\frac{k}{k_{\ast}}\right)}{\partial\nu_{s}}\Bigg|_{\nu_{s}=\frac{3}{2}}\right]$ (183)
@@ 7.2 Parity asymmetry versus even-odd asymmetry
$C_{\ell,\,BB}^{\rm SI}=\int_{0}^{\infty}\frac{dk}{k}A_{t}\left(\frac{k}{k_{\ast}}\right)^{n_{t}}j_{\ell}^{2}\left(\frac{k}{k_{s}}\right)$ (195)

### v2

@@ 6.2 The mathematical bridges in quantum black hole
Figure 5: The picture represents the spacetime conformal diagram of quantum SBH according to DQFT. It contains four regions I,II,III,IVI,\,II,\,III,\,IV, which define geometric SSS to describe quantum fields in Schwarzschild spacetime (applying the near-horizon approximation $r\approx 2GM$). In this picture, the regions II (IIIIII) and IIII (IVIV) are related by discrete transformation $\left(U,\,V\right)\to\left(-U,\,-V\right)$. The curved black lines with arrows represent integral curves of killing vector $\partial_{t}$ (54) in each region, or in other words, these are the curves of $T^{2}-X^{2}={\rm constant<0}$ (Region I and $I$) and $T^{2}-X^{2}={\rm constant}>0$ (Region III and IV). It is trivial to see an analogy between the spacetime of Schwarzschild BH and the phase space of IHO (23). The red horizontal lines are identified with $r=0$. The quantum field components in the exterior are $\hat{\Phi}_{I}$ and $\hat{\Phi}_{II}$ at parity conjugate regions with opposite arrows of time $t:$-$\infty\to\infty$ and $t:\infty\to$-$\infty$ respectively. Whereas in the interior quantum field (component) $\hat{\Phi}_{int}$ evolve according to (151).
@@ 6.2.1 A new understanding of ER=EPR
In all the investigations that admit unitarity is lost and usual entanglement between BH interior and exterior states, the basic fact that the concept of ”time” in the interior and exterior is different is unnoticed at the quantum level. It is well known that the radial coordinate $r$ behaves like time and $t$ becomes spacelike in the BH interior, which makes the interior Schwarzschild spacetime not static but analogous to the Kantowski-Sachs cosmological metric [97]. All QFTCS construction requires us to define a positive energy state with respect to an arrow of time, since time is not the same in the interior and exterior of BH; one must have a different Hilbert space structure for these two regions. This is exactly what our concept of geometric SSS does. Here, we present an alternative and complementary formulation of ER=EPR grounded in DQFT. By studying quantum fields across the Rindler and Schwarzschild horizons, we propose that the entanglement structure implied by ER=EPR arises naturally through the direct-sum Hilbert space formalism, without invoking geometric wormholes. To be precise, the outcomes of DQFT (137) and (154) indicate a pure state ($|\Psi_{P}\rangle$) is split into direct-sum of two pure state components $\left(|\Psi_{wH}\rangle,\,|\Psi_{bH}\rangle\right)$ that geometrically evolve within and beyond the Horizon as
@@ 6.3.1 Direct-sum QFT in flat FLRW de Sitter
$\hat{\phi}_{\pm}$ $=\int\frac{d^{3}k}{\left(2\pi\right)^{3/2}\sqrt{2k}}\left[d_{(\pm)\,\textbf{k}}\phi_{(\pm)k}\left(\pm\tau\right)e^{\pm i\textbf{k}\cdot\textbf{x}}+d^{\dagger}_{(\pm)\,\textbf{k}}\phi_{(\pm)k}^{\ast}\left(\pm\tau\right)e^{\mp i\textbf{k}\cdot\textbf{x}}\right]$
@@ 7 ERBs, direct-sum inflation (DSI) and CMB
This section marks a pivotal point in our exposition, where the theoretical framework developed thus far, including inverted harmonic oscillator dynamics, quantum field theory in spacetimes with horizons, and the direct-sum quantization formalism, culminates in potential observational consequences. Specifically, we propose a novel connection between the mathematical structures underpinning DQFT and anomalies observed in the Cosmic Microwave Background (CMB) [98], particularly those related to parity asymmetry. Furthermore, we provide a new prediction for primordial gravitational waves for the future observational probes.
The inflationary universe, modeled effectively by quasi-de Sitter spacetime, offers a natural setting where the horizon-induced quantum correlations may leave imprints on primordial fluctuations. In this context, we introduce the idea of Direct-Sum Inflation (DSI): a framework in which quantum states spanning parity conjugated regions, connected via mathematical bridges akin to those in ER=EPR, may manifest observable signatures in the large-scale CMB anisotropies. This opens a concrete avenue for testing the deep theoretical ideas developed throughout this work, providing a possible empirical handle on the quantum structure of spacetime near horizons.
@@ 7.1 Direct-sum QFT of inflationary quantum fluctuations
Inflationary background by definition quasi-dS expansion, and it breaks the symmetry (61) or (59) i.e., $\left(\tau,\,\textbf{x}\right)\to\left(-\tau,\,-\textbf{x}\right)$ by the time-dependent slow-roll parameters $\left(\epsilon,\,\eta\right)$ (82). Recall the fact that time is a parameter in quantum theory (in the sense that time reversal is anti-unitary in character), which we extensively discussed in Sec. 5. In the context of inflationary quantum fluctuations, we literally promote the gravitational and matter (inflaton) field variables as operators around the dynamically evolving quasi-de Sitter background. This is nothing but linearized quantum gravity, as extensively discussed in [99], because we do the following operationally (in the units of setting $8\pi G=1$).
@@
with $H^{(1)}_{\nu_{s}}\left(z\right)$ is the Hankel functions of the first kind. The standard inflation (SI) prediction [75] corresponds to $\Delta\mathcal{P}_{v}=0$, while the DSI one is given by Equation (177). Note that the power spectrum is the Fourier transform of the two-point function of curvature perturbation as shown in (176) and (177), which contains two unequal contributions due to $\mathcal{P}\mathcal{T}$ symmetry of dS (58) is broken during inflation. In other words our direct-sum vacuum $|0\rangle_{qdS}=|0_{+}\rangle_{qdS}\oplus|0_{-}\rangle_{qdS}$ is $\mathcal{P}\mathcal{T}$ asymmetric. We can express (178) as the two-point correlation of momentum modes as
@@
$|\Psi_{sq}\rangle=\begin{pmatrix}\prod_{\mathbf{k}}\frac{1}{\sqrt{|\alpha^{+}_{k}|}}\exp\!\Big(-\frac{\beta_{k}^{+}}{2\alpha_{k}^{+}}\,\hat{b}^{\dagger}_{\mathbf{k},\,+}\hat{b}_{-\textbf{k},\,+}^{\dagger}\Big)|0_{+}\rangle_{\rm qdS}\\ \prod_{\mathbf{k}}\frac{1}{\sqrt{|\alpha^{-}_{k}|}}\exp\!\Big(-\frac{\beta_{k}^{-}}{2\alpha_{k}^{-}}\,\hat{b}^{\dagger}_{\mathbf{k},\,-}\hat{b}_{-\textbf{k},\,-}^{\dagger}\Big)|0_{-}\rangle_{\rm qdS}\end{pmatrix}=\begin{pmatrix}|\Psi_{sq}^{+}\rangle\\ |\Psi_{sq}^{-}\rangle\end{pmatrix}$ (181)
@@
$P_{h}=A_{t}\left(\frac{k}{k_{\ast}}\right)^{n_{t}},\,r=-8n_{t}\,,\quad\Delta\mathcal{P}_{u}=\left(\frac{r}{8}\right)\operatorname{Re}\left[\frac{2}{H_{3/2}^{(1)}\left(\frac{k}{k_{\ast}}\right)}\frac{\partial H^{(1)}_{\nu_{s}}\left(\frac{k}{k_{\ast}}\right)}{\partial\nu_{s}}\Bigg|_{\nu_{s}=\frac{3}{2}}\right]$ (183)
@@ 7.2 Parity asymmetry versus even-odd asymmetry
$C_{\ell,\,BB}^{\rm SI}=\int_{0}^{\infty}\frac{dk}{k}A_{t}\left(\frac{k}{k_{\ast}}\right)^{n_{t}}T_{\ell,\,BB}^{2}\left(\frac{k}{k_{s}}\right)$ (195)